In a PTA fund raising activity, couples who came where charge P400.00. However if only one parent came, he was charged P250.00. 59 people attended the activity and P12,150.00 was raised. How many couples attended?

Answer by macston(5194)

.
C=number of couples; S=number of singles
.
2C+S=59
S=59-2C
.
P400C+P250S=P12150
P400C+P250(59-2C)=P12150
P400C+P14750-P500C=P12150
-P100C=-P2600
C=26
ANSWER 1: There were 26 couples.
.
S=59-2C
S=59-2(26)
S=59-52
S=7
ANSWER 2: There were 7 single parents.
.
CHECK:
P400C+P250S=P12150
P400(26)+P250(7)=P12150
P10400+P1750=P12150
P12150=P12150
.